Lourdes Corporation's 10% coupon rate, semiannual payment, $1,000 par value bonds, which mature in 10 years, are callable 4 years from today at $1,075. They sell at a price of $1,174.91, and the yield curve is flat. Assume that interest rates are expected to remain at their current level.

  1. What is the best estimate of these bonds' remaining life? Round your answer to the nearest whole number.
      years
  2. If Lourdes plans to raise additional capital and wants to use debt financing, what coupon rate would it have to set in order to issue new bonds at par?
    1. Since Lourdes wishes to issue new bonds at par value, the coupon rate set should be the same as that on the existing bonds.
    2. Since Lourdes wishes to issue new bonds at par value, the coupon rate set should be the same as the current yield on the existing bonds.
    3. Since interest rates have risen since the bond was first issued, the coupon rate should be set at a rate above the current coupon rate.
    4. Since the bonds are selling at a premium, the coupon rate should be set at the going rate, which is the YTC.
    5. Since the bonds are selling at a premium, the coupon rate should be set at the going rate, which is the YTM.
Answer #1

А В. 1 YTM 7.48% =RATE(10*2,100/2,-1174.91,1000)*2 2 YTc 6.62% =RATE(4* 2,100/2,-1174.91,1075)* 2 4 years, is remaining life

Please rate thumbs up